# CLiFT: Compressive Light-Field Tokens for Compute-Efficient and Adaptive Neural Rendering

### v1

@@ Abstract
This paper proposes a neural rendering approach that represents a scene as “compressed light-field tokens (CLiFTs)”, retaining rich appearance and geometric information of a scene. CLiFT enables compute-efficient rendering by compressed tokens, while being capable of changing the number of tokens to represent a scene or render a novel view with one trained network. Concretely, given a set of images, multi-view encoder tokenizes the images with the camera poses. Latent-space K-means selects a reduced set of rays as cluster centroids using the tokens. The multi-view “condenser” compresses the information of all the tokens into the centroid tokens to construct CLiFTs. At test time, given a target view and a compute budget (i.e., the number of CLiFTs), the system collects the specified number of nearby tokens and synthesizes a novel view using a compute-adaptive renderer. Extensive experiments on RealEstate10K and DL3DV datasets quantitatively and qualitatively validate our approach, achieving significant data reduction with comparable rendering quality and the highest overall rendering score, while providing trade-offs of data size, rendering quality, and rendering speed. Check demos and code on the project page: https://c-lift.github.io.
@@ 3 Compressive Light-Field Tokens (CLiFT)
Figure 1: The training and the inference system overview. Top: The training consists of three steps: 1) Multi-view encoder, tokenizing the input images; 2) Latent K-means, selecting a representative set of tokens; and 3) Neural condensation, compressing the information of all the tokens into the representative set to produce Compressed Light-Field Tokens (CLiFTs). Bottom: At inference time, given a target view, we collect a relevant set of rays with CLiFTs and render a novel view.
A light field is a collection of half-rays, each associated with the radiance information. Compressive light-field tokens (CLiFTs) are a collection of half-rays, each associated with a latent vector to which a neural encoder compresses the geometry and radiance information of a scene (See Figure 1). Given a target camera pose, a traditional light field rendering synthesizes a color for each target ray, typically by linear interpolation of nearby input rays. CLiFT rendering synthesizes an image (not necessarily per ray) by using nearby CLiFTs via a neural renderer. The framework allows us to control data size, rendering quality, and rendering speed by two hyper parameters 1) Storage CLiFT count ($N_{s}$), the number of tokens to represent a scene; and 2) Render CLiFT count ($N_{r}$), the number of tokens to use in rendering a frame. The section explains the problem, CLiFT construction, and CLiFT rendering.
@@ 3.1 Problem Definition
This paper tackles the problem of taking $N_{c}$ images with the camera poses, and constructing $N_{s}$ CLiFTs as a scene representation. Given a target view, the output is its synthesized image, where $N_{r}$ CLiFTs tokens are to be selected and used for rendering. The evaluation is based on standard image reconstruction metrics, PSNR, SSIM, and LPIPS. A token (CLiFT) is a $D$ dimensional embedding, associated with a half-ray. Ray geometries are represented in a global coordinate frame.
@@ 3.3 CLiFT Rendering
Token Selection. The token selection algorithm employs a simple heuristic. Let $N_{r}$ denote the number of CLiFTs used for rendering. To ensure spatial coverage of the target view, we divide the view into a 16$\times$16 grid of patches and pad it with a 4-patch margin on all sides, resulting in a 24$\times$24 grid. For each patch, we cast a ray through its center and retrieve the $N_{r}/(24\times 24)$ closest CLiFTs from a pool of $N_{s}$ storage CLiFTs. The distance between a patch and a CLiFT is computed using a heuristic based on their associated half-rays (ray origins and directions); details are provided in the supplementary material. Since CLiFTs may be selected multiple times, we greedily accumulate CLiFTs in order of their minimum distance to any patch ray until $N_{r}$ unique tokens are selected.
@@ 4.2 Main Results
Figure 2 and Figure 4 show the main evaluation results on the RealEstate10K and DL3DV datasets. In the plots, the x-axis represents the data size of the scene representation: Storage CLiFTs for our method, decoder input tokens for LVSM, and splats for MVSplat and DepthSplat. The y-axis reports PSNR; due to space constraints, the LPIPS and SSIM plots are deferred to the Appendix, where they show consistent trends.
@@ Acknowledgments
Please check our project page at https://c-lift.github.io, which shows rendering results at different compression rates and compares with the baseline methods.

### v2

# CLiFT: Compressive Light-Field Tokens for Compute Efficient and Adaptive Neural Rendering

@@ Abstract
This paper proposes a neural rendering approach that represents a scene as “compressed light-field tokens (CLiFTs)”, retaining rich appearance and geometric information of a scene. CLiFT enables compute-efficient rendering by compressed tokens, while being capable of changing the number of tokens to represent a scene or render a novel view with one trained network. Concretely, given a set of images, multi-view encoder tokenizes the images with the camera poses. Latent-space K-means selects a reduced set of rays as cluster centroids using the tokens. The multi-view “condenser” compresses the information of all the tokens into the centroid tokens to construct CLiFTs. At test time, given a target view and a compute budget (i.e., the number of CLiFTs), the system collects the specified number of nearby tokens and synthesizes a novel view using a compute-adaptive renderer. Extensive experiments on RealEstate10K and DL3DV datasets quantitatively and qualitatively validate our approach, achieving significant data reduction with comparable rendering quality and the highest overall rendering score, while providing trade-offs of data size, rendering quality, and rendering speed. Check demos and code on the project page: https://clift-nvs.github.io.
@@ 3 Compressive Light-Field Tokens (CLiFT)
Figure 1: The training and the inference system overview. Top: The training consists of three steps: 1) Multi-view encoder, tokenizing the input images; 2) Latent K-means, selecting a representative set of tokens; and 3) Neural condensation, compressing the information of all the tokens into the representative set to produce Compressed Light-Field Tokens (CLiFTs). Bottom: At inference time, multi-view images are encoded into CLiFTs following the same process as in training. Given a target view, we collect a relevant set of CLiFTs with simple heuristics and render a novel view.
A light field is a collection of rays, each associated with the radiance information. Compressive light-field tokens (CLiFTs) are a collection of rays, each associated with a latent vector to which a neural encoder compresses the geometry and radiance information of a scene (See Figure 1). Given a target camera pose, a traditional light field rendering synthesizes a color for each target ray, typically by linear interpolation of nearby input rays. CLiFT rendering synthesizes an image (not necessarily per ray) by using nearby CLiFTs via a neural renderer. The framework allows us to control data size, rendering quality, and rendering speed by two hyper parameters 1) Storage CLiFT count ($N_{s}$), the number of tokens to represent a scene; and 2) Render CLiFT count ($N_{r}$), the number of tokens to use in rendering a frame. The section explains the problem, CLiFT construction, and CLiFT rendering.
@@ 3.1 Problem Definition
This paper tackles the problem of taking $N_{c}$ images with the camera poses, and constructing $N_{s}$ CLiFTs as a scene representation. Given a target view, the output is its synthesized image, where $N_{r}$ CLiFTs tokens are to be selected and used for rendering. The evaluation is based on standard image reconstruction metrics, PSNR, SSIM, and LPIPS. A token (CLiFT) is a $D$ dimensional embedding, associated with a ray. Ray geometries are represented in a global coordinate frame.
@@ 3.3 CLiFT Rendering
Token Selection. The token selection algorithm employs a simple heuristic. Let $N_{r}$ denote the number of CLiFTs used for rendering. To ensure spatial coverage of the target view, we divide the view into a 16$\times$16 grid of patches and pad it with a 4-patch margin on all sides, resulting in a 24$\times$24 grid. For each patch, we cast a ray through its center and retrieve the $N_{r}/(24\times 24)$ closest CLiFTs from a pool of $N_{s}$ storage CLiFTs. The distance between a patch and a CLiFT is computed using a heuristic based on their associated rays (ray origins and directions); details are provided in the supplementary material. Since CLiFTs may be selected multiple times, we greedily accumulate CLiFTs in order of their minimum distance to any patch ray until $N_{r}$ unique tokens are selected.
@@ 4.2 Main Results
Figure 2 and Figure 4 show the main evaluation results on the RealEstate10K and DL3DV datasets. In the plots, the x-axis represents the data size of the scene representation: Storage CLiFTs for our method, decoder input tokens for LVSM, and splats for MVSplat and DepthSplat. The y-axis reports PSNR; due to space constraints, the LPIPS and SSIM plots are deferred to the supplementary material, where they show consistent trends.
@@ Acknowledgments
Please check our project page at https://clift-nvs.github.io, which shows rendering results at different compression rates and compares with the baseline methods.